a solid body starts moving about a stationery axis with an angular accelration '0.02t'rad/s here 't' is in sec. How soon after the beginning of rotation, will the accelartion vector of an arbitary point of the body form an angle 60 degree with the velocity vector at the same point

Grade:12

1 Answers

vikas askiitian expert
13 years ago

angular accleration =0.02t

 ler r be distance from axis of rotation...

tangential accleration =0.02tr

velocity at any time =adt=0.01rt^2

total accleration(A) =sqrt ( (ac)^2 +(at)^2 )

ac and at are centripital and tangential accleration ....

total accleration makes 60 degree angle with velocity or with tangential accleration....

Acos60=(at)

after solving we get

ac/at=sqrt3

ac=v^2/r and at=0.02tr

t=7 sec ans